# Jets-separation thresholds, Seshadri constants and higher Gauss-Wahl maps on abelian varieties Nelson Alvarado N. Alvarado

### v1

@@ 6. Effective surjectivity of Gauss-Wahl maps
The surjectivity of (6.1) follows as soon Indeed: this inequality means that the -twisted sheaf is IT(0), where is the class of and in which in particular means that and thus that is surjective.
@@
Let be an abelian variety and an ample line bundle on it. Fix a rational number with Then for every with and for every divisible enough integer number, there exist a divisor such that
@@
where is the class of and Now, as the above inequality becomes (6.2). The inequality of follows as in the proof of Theorem 6.1.
@@ 7. Further questions
2. The theory of the Fourier-Mukai functors, generic vanishing and cohomological rank functions work also over algebraically closed fields of positive characteristic. In this context, most of the results of this work also hold in that setting. However, in positive characteristic Nadel’s vanishing theorem does not hold and hence we are not able to prove Proposition 6.3. We may then ask the following:
@@
Let be a polarized abelian variety over a field of positive characteristic. Does the inequality hold?

### v2

@@ 6. Effective surjectivity of Gauss-Wahl maps
The surjectivity of (6.1) follows as soon as Indeed, by Lemma 5.3b) this inequality means that the -twisted sheaf is IT(0), where is the class of and in which in particular means that and thus that is surjective.
@@
Let be an abelian variety and an ample line bundle on it. Fix a rational number with Then for every with and for every divisible enough integer, there exists a divisor such that
@@
where is the class of and Now, as the above inequality becomes (6.2).
@@ 7. Further questions
2. Note that, except for Proposition 6.3, whose proof uses Nadel’s vanishing theorem, the results presented in this article hold over an algebraically closed field of any characteristic. We may then ask the following:
@@
Let be a polarized abelian variety over a field of positive characteristic. Does the inequality still hold?
@@
This work is part of the author’s Phd thesis. The author would like to thank his advisor, Giuseppe Pareschi, for his invaluable guidance and help to improve both the exposition and the mathematical content of this article. The author also thanks to the anonymous referees for their comments, which were very useful to improve the presentation of this article and to clarify some proofs. This work was supported by the MIUR Excellence Department Project MatMod@TOV awarded to the Department of Mathematics of the University of Rome Tor Vergata.